Grade 12th passMechanics

A bullet is fired horizontally on a150m away target with speed 1500m/sec.How much below does the bullet hit the target?

nilanjana
8 Years agoGrade 12th pass

1 Answer

Nevin
8 Years ago
The velocity of the bullet in 1500 m/s
The target is at a distance of 150 m
So the bullet takes 0.1 s to reach the target
The initial velocity along vertical component is 0 m/s
We know that s = ut + 1/2 at square 
Therefore s = 0 + 1/2 (9.8)*(0.1) square
s = 4.9*0.01
s = 0.049 m
So the bullet hit 0.049 m below the target